# EXIT: Context-Aware Extractive Compression for Enhancing Retrieval-Augmented Generation

### v2

@@ B.2 Performance with Proprietary Model
B.3 Impact of Threshold $\tau$
Figure 7: Changes in EM, F1 score, and token count as the threshold $T$ for retaining sentences is adjusted.
@@ B.3 Impact of Threshold $\tau$
B.4 Analysis of Classification Performance Across Negative Sample Types
Table 8: Classification performance for Yes/No labels, including overall accuracy.
@@ B.4 Analysis of Classification Performance Across Negative Sample Types
B.5 Classification Performance under Ablation Settings
B.5.1 Analysis of Training Data Composition
Figure 9 presents row-normalized confusion matrices comparing classification performance across three training data configurations: Ours (Pos+H-Neg+Neg), Pos+H-Neg, and Pos+Neg. Under the Ours setup, the classifier displays a balanced ability to identify both “Yes” (relevant) and “No” (irrelevant) sentences, achieving an F1-score of 0.92 for both classes. In contrast, excluding one type of negative sample (Pos+H-Neg or Pos+Neg) reduces overall robustness, evidenced by declines in both accuracy and class-wise F1 scores. For instance, the Pos+Neg configuration struggles to maintain balance, accurately identifying “Yes” instances but misclassifying a substantial number of “No” cases. These results confirm that incorporating a comprehensive mix of positive, hard-negative, and random-negative samples leads to more reliable and contextually aware sentence selection, thereby improving the classifier’s performance in practical retrieval-augmented QA scenarios.
B.5.2 Impact of Context on Classification Performance
@@
B.6 Training Data Ablation Analysis
Table 9 compares models trained on HQA, 2WIKI, or both. Training solely on HQA yields the highest EM and F1 (31.6 EM, 42.6 F1) with moderate token usage. In contrast, 2WIKI training improves compression but lowers accuracy (29.2 EM, 40.3 F1). Combining datasets does not surpass HQA alone.
@@ B.6 Training Data Ablation Analysis
B.7 Case Studies
@@
In Table 10, the original documents contain the correct answer (“Custard”) but also include distracting information (“Eggnog”). Despite having the necessary evidence, the reader fails to produce the correct answer, likely due to this distractor. In contrast, our method (Ours) filters out irrelevant details, drastically reduces input length, and retains only the essential context needed to answer the query accurately. As a result, the reader confidently generates the correct answer (“Custard”).
In a second scenario (Table 11), the original documents retrieve multiple documents related to “Wagner” or “sci-fi” series but fail to provide any content explicitly linking James Belushi to the correct 90s sci-fi series, resulting in an incorrect prediction. Surprisingly, Ours removes all retrieved context entirely, providing the reader with no additional information. Under this no-context condition, the reader relies solely on its internal knowledge and, in this case, correctly identifies “Wild Palms.” While this outcome indicates a form of hallucination or model bias—since the answer emerges without external supporting evidence—it also demonstrates Ours’ capability to avoid misleading context. By eliminating irrelevant or confusing documents, Ours can sometimes allow the model’s internal knowledge to surface, leading to correct answers even in the absence of any retrieved information.
@@
Table 12 compares Ours with several competing compression approaches. CompAct preserves some relevant information but introduces hallucinations, causing the reader to claim that it cannot find the correct answer. Refiner omits the crucial entity required to answer the query, demonstrating how abstractive compressors may inadvertently remove key content. In contrast, Ours avoids hallucinations and retains the answer’s entity in a concise, coherent form.
LongLLMLingua’s token-level filtering approach yields unreadable text and removes the essential “Romania” entity, preventing the reader from generating the correct answer. Ours, on the other hand, maintains semantic coherence and includes the correct entity, allowing the reader to produce the correct answer without interference.
These case studies highlight the advantages of Ours: it eliminates distractors, preserves critical entities, and maintains semantic integrity. Consequently, the reader consistently arrives at correct answers with fewer tokens and no hallucinations.

### v3

@@ C.2 Performance with Proprietary Model
C.3 Evaluation in Specialized Domains
Table 10: Evaluation of EXIT and baselines in specialized domains. BioASQ evaluates biomedical QA using PubMed documents, and COVID-QA focuses on COVID-specific QA over the TREC-COVID corpus.
BioASQ (Biomedical QA)
Method EM F1 Avg. #Tokens
Original Docs 57.2 68.6 1791.49
RECOMP-Abst 54.6 63.8 37.72
CompAct 58.0 68.2 94.81
Refiner 58.8 68.5 229.22
RECOMP-Extr 55.8 65.0 109.72
LongLLMLingua 58.2 66.4 364.06
EXIT (Ours) 58.6 67.1 132.28
COVID-QA (COVID-19 QA)
Method EM F1 Avg. #Tokens
Original Docs 3.8 16.0 1775.66
RECOMP-Abst 2.2 11.5 17.94
CompAct 3.4 14.6 95.65
Refiner 3.8 14.6 201.38
RECOMP-Extr 1.4 11.4 103.11
LongLLMLingua 2.2 13.5 353.03
EXIT (Ours) 5.8 17.2 129.33
To assess the generalizability of EXIT beyond general-domain datasets, we conducted additional evaluations on two specialized QA benchmarks: 1) BioASQ Tsatsaronis et al. (2015), a biomedical QA dataset using the PubMed corpus (14.9M documents), 2) COVID-QA Möller et al. (2020), a COVID-specific QA benchmark based on the TREC-COVID collection Roberts et al. (2021) (171K documents).
Although our classifier was trained solely on HQA, a general-domain dataset, EXIT performs competitively or better than existing baselines in both specialized domains. Table 10 shows that EXIT achieves high EM and F1 scores while substantially reducing token counts. On COVID-QA, EXIT notably outperforms all baselines in EM and F1 despite the domain shift.
These results suggest that EXIT’s context-aware extraction strategy generalizes well, preserving key evidence without requiring domain-specific fine-tuning. This highlights EXIT’s robustness across diverse retrieval scenarios.
C.4 Evaluation on Complex Multi-hop Reasoning
Table 11: Evaluation of EXIT and baselines on the MuSiQue dataset on Llama3-8B reader.
Method EM F1 Latency (s)
Original Docs 3.8 10.3 1.1
RECOMP-Abst 4.5 11.8 2.5
CompAct 4.1 11.0 8.5
Refiner 4.5 11.3 7.6
RECOMP-Extr 4.3 10.1 0.6
LongLLMLingua 4.0 10.7 0.9
EXIT (Ours) 4.6 11.3 0.9
To further assess EXIT’s ability to handle complex reasoning tasks, we evaluated it on the MuSiQue dataset Trivedi et al. (2022), which contains multi-hop questions requiring deeper reasoning (3–4 hops). While our main experiments focused on standard benchmarks like HQA and 2WIKI, MuSiQue provides a valuable test of generalization under more challenging conditions.
As shown in Table 11, EXIT achieves competitive performance on MuSiQue. Notably, it maintains low latency while matching or outperforming other extractive and abstractive baselines in EM and F1 scores. These results demonstrate EXIT’s robustness in handling longer reasoning chains without incurring the high latency costs typically associated with abstractive methods.
These findings reinforce EXIT’s ability to scale to more complex, multi-hop QA settings while maintaining an efficient trade-off between latency and performance.
C.5 Latency Measurement and Analysis
Table 12: Average end-to-end latency (seconds per sample) across different compression methods. Results are the mean and standard deviation (Std) over five runs. Experiments for the 8B reader were conducted on a single A100-80GB SXM GPU, and for the 70B reader on four A100-80GB SXM GPUs.
8B Reader Compression Reading Total
Method Avg Std Avg Std Avg Std
Original Docs 0.9481 0.0056 0.9481 0.0056
RECOMP-Abst 1.9967 0.0159 0.3177 0.0008 2.3144 0.0167
CompAct 12.4682 0.0030 0.3930 0.0006 12.8612 0.0036
Refiner 6.3881 0.0332 0.3877 0.0012 6.7758 0.0344
RECOMP-Extr 0.0279 0.0022 0.3668 0.0040 0.3947 0.0062
LongLLMLingua 0.3721 0.0050 0.4526 0.0004 0.8247 0.0054
EXIT (Ours) 0.4263 0.0031 0.4529 0.0028 0.8792 0.0059
70B Reader Compression Reading Total
Method Avg Std Avg Std Avg Std
Original Docs 8.0835 0.0072 8.0835 0.0072
RECOMP-Abst 1.8621 0.0171 2.3210 0.0113 4.1831 0.0284
CompAct 15.1999 0.0132 2.6932 0.0007 17.8931 0.0139
Refiner 8.3520 0.0248 3.0993 0.0017 11.4513 0.0265
RECOMP-Extr 0.0329 0.0069 2.8288 0.0031 2.8617 0.0100
LongLLMLingua 0.4788 0.0127 3.7121 0.0012 4.1909 0.0139
EXIT (Ours) 0.3253 0.0086 3.5341 0.0022 3.8594 0.0108
To ensure the reliability and reproducibility of our latency results, we conducted five independent runs for each compression method and report both the mean and standard deviation of end-to-end latency (measured in seconds per sample). End-to-end latency includes both compression time and reading/inference time, measured from the moment a request is issued to the moment the answer is returned by the model.
For consistency, all experiments using the 8B reader were run on a single A100-80GB SXM GPU, while those using the 70B reader were executed on four A100-80GB SXM GPUs. To minimize hardware-induced noise, we ensured that no unrelated processes were active during measurement and maintained fixed hyperparameters and model configurations across all runs.
As shown in Table 12, our method EXIT achieves a strong balance between latency and accuracy. While abstractive methods such as CompAct and Refiner incur significant latency (12.86s and 6.78s ,respectively, with the 8B reader), EXIT completes inference in under 1 second (0.88s) on average. Even with the 70B reader, EXIT remains efficient with 3.86s total latency, significantly faster than the original uncompressed baseline (8.08s), while also improving QA performance.
These results reinforce EXIT’s practicality for real-world deployment, particularly in latency-sensitive applications.
C.6 Impact of Threshold $\tau$
Figure 7: Changes in EM, F1 score, and token count as the threshold $\tau$ for retaining sentences is adjusted.
@@ C.6 Impact of Threshold $\tau$
C.7 Effect of Classifier Training
Table 13: Comparison of EXIT using a zero-shot (frozen Gemma-2B) vs. fine-tuned classifier on HotpotQA (HQA) and 2WikiMultiHopQA (2WIKI).
Dataset Method EM F1
2WIKI Original Docs 18.0 25.7
EXIT (Zero-shot) 22.4 27.9
EXIT (Ours) 24.8 30.1
HQA Original Docs 29.2 40.2
EXIT (Zero-shot) 31.4 43.1
EXIT (Ours) 31.6 42.6
Dataset Method Avg. #Tokens
2WIKI Original Docs 764.47
EXIT (Zero-shot) 731.35
EXIT (Ours) 145.62
HQA Original Docs 735.24
EXIT (Zero-shot) 727.87
EXIT (Ours) 191.11
To evaluate the impact of fine-tuning the classifier in EXIT, we compared our trained model with a frozen zero-shot version using Gemma-2B (without any task-specific fine-tuning). We tested both variants on an in-domain dataset (HQA) and an out-of-domain dataset (2WIKI).
As shown in Table 13, the zero-shot classifier already improves over the uncompressed baseline, indicating the strength of relevance-based filtering even without training. However, our fine-tuned classifier consistently outperforms the zero-shot variant in both EM and F1 scores while achieving significantly greater token reduction—demonstrating that supervised training yields both better accuracy and more efficient compression.
C.8 Adaptability of EXIT to Different Supervision Signals
Table 14: Comparison of EXIT with baselines on QA tasks on HQA. EXIT (GPT-4o) refers to a variant using GPT-4o as a zero-shot relevance classifier, while EXIT (Ours) represents our proposed method trained with supervised data.
Method EM F1 #token
Original Docs 29.2 40.2 735.2
RECOMP-Extr 25.2 34.8 89.7
RECOMP-Abst 19.8 24.7 55.5
LongLLMLingua 27.4 40.2 227.1
CompAct 29.4 40.9 76.3
Refiner 28.8 40.7 73.4
EXIT (GPT-4o) 30.4 41.9 84.2
EXIT (Ours) 31.6 42.6 191.0
A key strength of EXIT is its ability to integrate different classifiers for sentence selection without being constrained to manually annotated datasets. To explore this flexibility, we evaluate an alternative approach where relevance scores are derived from GPT-4o without explicit fine-tuning. As shown in Table 14, EXIT maintains strong performance, outperforming or closely matching baselines such as LongLLMLingua and CompAct. This demonstrates that EXIT can leverage diverse supervision signals while remaining adaptable to different scoring mechanisms. Moreover, these results suggest the potential of utilizing large-scale pseudo-labeled data to further refine EXIT’s training, enhancing scalability without relying strictly on human-labeled datasets. This adaptability highlights EXIT’s robustness and practical applicability across various retrieval settings.
C.9 Analysis of Classification Performance Across Negative Sample Types
Table 15: Classification performance for Yes/No labels.
@@ C.9 Analysis of Classification Performance Across Negative Sample Types
C.10 Classification Performance
To better understand the effectiveness of our context-aware relevance classifier, we report row-normalized confusion matrices for both in-domain (HQA) and out-of-domain (2WIKI) datasets, as shown in Figure 8. On HQA, the classifier displays a perfectly balanced ability to recognize both relevant (“Yes”) and irrelevant (“No”) sentences, achieving over 90% precision and recall in each category. While, on the 2WIKI dataset, the classifier exhibits a slight drop in recall for “Yes” sentences, it still shows a strong classification ability with over 70% recall and 90% precision.
These results confirm that the classifier performs robustly in its training domain and generalizes reasonably well to unseen queries, though we leave narrowing this discrepancy as a valuable future research direction.
C.11 Classification Performance under Ablation Settings
C.11.1 Analysis of Training Data Composition
Figure 9 presents row-normalized confusion matrices comparing classification performance across three training data configurations: Ours (Pos+H-Neg+Neg), Pos+H-Neg, and Pos+Neg. Under the Ours setup, the classifier displays a balanced ability to identify both “Yes” (relevant) and “No” (irrelevant) sentences, achieving an F1-score of 0.92 for both classes. By contrast, excluding one type of negative sample (Pos+H-Neg or Pos+Neg) reduces overall robustness, evidenced by declines in both accuracy and class-wise F1 scores. For instance, the Pos+Neg configuration struggles to maintain balance, accurately identifying “Yes” instances but misclassifying a substantial number of “No” cases. These results confirm that incorporating a comprehensive mix of positive, hard-negative, and random-negative samples leads to more reliable and contextually aware sentence selection, thereby improving the classifier’s performance in practical retrieval-augmented QA scenarios.
C.11.2 Impact of Context on Classification Performance
@@
C.12 Sentence Reordering Strategy Analysis
Table 16: Ablation results comparing sentence reordering strategies in EXIT on HotpotQA. “Random” shuffles selected sentences, while “Ascending” and “Descending” sort by relevance scores. “Original order” maintains the sentence sequence from the original document.
Reordering Strategy EM F1
Random 30.20 41.30
Ascending (score-based) 30.43 41.44
Descending (score-based) 30.33 41.16
Original order (EXIT) 30.61 41.49
One potential concern with extractive compression is that removing many sentences may lead to unnatural or ambiguous contexts when the remaining content is reassembled. In EXIT, we preserve the original sentence order during reconstruction to maintain document coherence, following prior work Hwang et al. (2024). However, we acknowledge that this decision could lead to suboptimal flow when large gaps exist between selected sentences.
To assess this design choice, we conducted an ablation study comparing four sentence reordering strategies: 1) Random: randomly shuffles selected sentences, 2) Ascending: sorts by relevance score (low to high), 3) Descending: sorts by relevance score (high to low), 4) Original order (EXIT): preserves sentence sequence from the source document
Table 16 shows that maintaining the original order consistently yields the best EM and F1 scores on HQA. This confirms that preserving the source structure helps retain contextual flow, even when intermediate content is omitted. We include this result to support our design and to acknowledge the known limitations of extractive methods in handling discourse coherence.
C.13 Training Data Ablation Analysis
@@
Table 17 compares models trained on HQA, 2WIKI, or both. Training solely on HQA yields the highest EM and F1 (31.6 EM, 42.6 F1) with moderate token usage. By contrast, 2WIKI training improves compression but lowers accuracy (29.2 EM, 40.3 F1). Combining datasets does not surpass HQA alone.
C.14 Case Studies
@@
In Table 18, the original documents contain the correct answer (“Custard”) but also include distracting information (“Eggnog”). Despite having the necessary evidence, the reader fails to produce the correct answer, likely due to this distractor. By contrast, EXIT (Ours) filters out irrelevant details, drastically reduces input length, and retains only the essential context needed to answer the query accurately. As a result, the reader confidently generates the correct answer (“Custard”).
In a second scenario (Table 19), the original documents retrieve multiple documents related to “Wagner” or “sci-fi” series but fail to provide any content explicitly linking James Belushi to the correct 90s sci-fi series, resulting in an incorrect prediction. Surprisingly, EXIT removes all retrieved context entirely, providing the reader with no additional information. Under this no-context condition, the reader relies solely on its internal knowledge and, in this case, correctly identifies “Wild Palms.” While this outcome indicates a form of hallucination or model bias—since the answer emerges without external supporting evidence—it also demonstrates EXIT’s capability to avoid misleading context. By eliminating irrelevant or confusing documents, EXIT can sometimes allow the model’s internal knowledge to surface, leading to correct answers even in the absence of any retrieved information.
@@
Table 20 compares EXIT with several competing compression approaches. CompAct preserves some relevant information but introduces hallucinations, causing the reader to claim that it cannot find the correct answer. Refiner omits the crucial entity required to answer the query, demonstrating how abstractive compressors may inadvertently remove key content. By contrast, EXIT avoids hallucinations and retains the answer’s entity in a concise, coherent form.
LongLLMLingua’s token-level filtering approach yields unreadable text and removes the essential “Romania” entity, preventing the reader from generating the correct answer. EXIT, on the other hand, maintains semantic coherence and includes the correct entity, allowing the reader to produce the correct answer without interference.
These case studies highlight the advantages of EXIT: it eliminates distractors, preserves critical entities, and maintains semantic integrity. Consequently, the reader consistently arrives at correct answers with fewer tokens and no hallucinations.